OpenStudy (anonymous):

Part 1: Is the following equation true (2 points) Part 2: Part 2: Use complete sentences to explain the properties in making your decision. (6 points) \[4 \log _{3} x + \frac{ 1 }{ 2 } \log _{3} y - 2 \log _{3} z =\log \frac{ x ^{4}\sqrt{y} }{ z ^{2} }\]

OpenStudy (campbell_st):

well this uses 3 different log laws 1. powers \[a \log(b) = \log(b^a)\] 2. addition & multiplication \[\log(a) + \log(b) = \log(ab)\] 3. subtraction/division \[\log(a) - \log(b) = \log(\frac{a}{b})\] you need to use all 3 start by using the power law for all 3 logs...